OpenStudy (anonymous):

Gregory sees an $80.00 jacket on sale at 30% off. How much will it cost after a 7% sales tax?

OpenStudy (mathmate):

Sale price = list price * (1-discount) Final price with tax = sale price * (1+ tax rate) Example, A car listed at $20000 is offered at a discount of 10%. What is the amount payable with a sales tax of 9%. Sale price = 20000*(1-10%)=20000*0.9=18000 Amount payable = 18000*(1+0.09)=18000*1.09=19620 Answer: Amount payable is $19620.

OpenStudy (anonymous):

ok thanks

OpenStudy (mathmate):

You're welcome! :)